# SeisFusion: Constrained Diffusion Model with Input Guidance for 3D Seismic Data Interpolation and Reconstruction

### v1

@@ IV-D2 Mobil Avo Viking Graben Line 12
Random discrete missing: we set two different missing rates, 50% and 80%, where traces were set to 0 to represent the missing data. Fig. 9 displays the reconstruction results of different methods under these two missing rates. It can be observed that the reconstruction results of UNet and the original diffusion model already show low consistency at 50% missing rate, and as the missing rate increases, the inconsistency becomes more apparent, as indicated by the residual plots showing significant deviations from the ground truth. Both MDA GAN and our method performed well in reconstructing under both missing scenarios, with our method slightly outperforming MDA GAN according to the residual plots.
@@
Random Missing 50% Traces UNet 5.1437e-3 29.3055 0.9173
Diffusion Model 5.6712e-3 22.9027 0.8482
@@
Random Missing 80% Traces Missing 9.7737e-3 17.5117 0.8597
Traces 9.8717e-3 21.5047 0.8061
@@
Continuous missing: we similarly set two scenarios: continuous missing of 50 traces and continuous missing of 100 traces. Traces were set to 0 to represent the missing data, and the reconstruction results of the four models are shown in Fig. 10. It can be observed that UNet fails to reconstruct seismic data under continuous missing, and the original diffusion model, due to lack of constraints and sampling guidance, produces discontinuous and inconsistent data under 50-trace continuous missing, which becomes more pronounced with 100-trace continuous missing. According to the residual plots, significant deviations from the ground truth are observed, particularly at larger missing values. MDA GAN performs relatively well in reconstruction, but deviations and inconsistencies still appear at the missing boundaries. Our method provides optimal reconstruction results under different missing scenarios.
@@
50 continuous missing traces UNet 0.0505 12.9651 0.7792
Diffusion Model 6.1376e-3 21.9251 0.8418
@@
100 continuous missing traces Missing 0.1012 9.9480 0.5788
Traces 0.0987 21.6757 0.7321
@@
V Conclusions

### v2

@@ IV-D2 Mobil Avo Viking Graben Line 12
Random discrete missing: we set two different missing rates, 50% and 80%, where traces were set to 0 to represent the missing data. Fig. 9 displays the reconstruction results of different methods under these two missing rates. It can be observed that the reconstruction results of UNet and the conditional DPM already show low consistency at 50% missing rate, and as the missing rate increases, the inconsistency becomes more apparent, as indicated by the residual plots showing significant deviations from the ground truth. Both MDA GAN and our method performed well in reconstructing under both missing scenarios, with our method slightly outperforming MDA GAN according to the residual plots.
@@
Random Missing 50% Traces UNet 7.4965e-4 31.2513 0.9358
Conditional DPM 1.0068e-3 29.9702 0.9213
@@
Random Missing 80% Traces UNet 1.6030e-3 27.9505 0.8987
Conditional DPM 3.6524e-3 26.5859 0.8947
@@
Continuous missing: we similarly set two scenarios: continuous missing of 50 traces and continuous missing of 100 traces. Traces were set to 0 to represent the missing data, and the reconstruction results of the four models are shown in Fig. 10. It can be observed that UNet fails to reconstruct seismic data under continuous missing, reconstructing undesired data in the low-amplitude part, and the conditional DPM reconstructed data show some inconsistency with high values overall. MDA GAN performs relatively well in reconstruction, but deviations and inconsistencies still appear at the missing boundaries. Our method provides optimal reconstruction results under different missing scenarios.
@@
50 continuous missing traces UNet 4.6285e-4 33.3455 0.9617
Conditional DPM 3.0986e-4 34.5915 0.9649
@@
100 continuous missing traces UNet 1.3854e-3 28.5839 0.9015
Conditional DPM 1.9679e-3 29.0143 0.9263
@@
Figure 11: (a) Ground truth,(b) 50 continuous missing traces, (a-1) to (a-4) are the reconstruction results of Unet, conditional DPM, MDA GAN, and SeisFusion, respectively. (b-1) to (b-4) are the residuals of each reconstruction result relative to the ground truth.
V Discussion
V-A Generalization
To explore the generalization ability of the proposed method, we conducted generalization experiments. The training set for the experiment consisted of synthetic data SEG C3. The model trained on SEG C3 was then directly used to reconstruct a more complex field dataset Mobil Avo Viking Graben Line 12. We intentionally removed 50 consecutive traces and set the traces to zero to represent the missing traces, which was used as the data to be reconstructed. The reconstruction results of the four models are shown in Fig. 11. It can be observed that the generalization ability of convolution-based reconstruction methods, Unet and MDA GAN, is relatively weak. In the low amplitude portion of the upper triangle in the reconstructed data, they produced features similar to those of the SEG C3 dataset. This may be because convolution-based methods are a point-to-point reconstruction approach, learning the characteristics of the training set. When applied to untrained data with significant differences, this point-to-point reconstruction method can only mechanically reconstruct based on the characteristics of the training set. In contrast, the diffusion model-based reconstruction methods, DPM and our method, were able to reconstruct the basic features of the data. In the low amplitude portion of the upper triangle, the reconstructed data showed high consistency. This is largely due to the fact that the diffusion models learn the probability distribution of the target data and can provide an ideal coverage of the distribution, thus having stronger generalization ability. However, in the high amplitude portion of the lower triangle, there were more noticeable inconsistencies.
Figure 12: Feature maps extracted by the network from SEG C3 data with 50% random missing at times t=5, t=50, and t=200, along with estimates for $\epsilon_{\theta}(x_{t},y,t)$ and sampled data.
We calculated four evaluation metrics in Table VI to quantitatively assess the generalization ability of the four methods. The results show that the methods based on diffusion models have strong generalization ability. Moreover, thanks to guided sampling, our method outperforms DPM in terms of generalization ability and is already capable of preliminarily reconstructing more complex data.
TABLE VI: Comparison of Generalization Ability of Four Reconstruction Networks with 50 Consecutive Missing Traces in Field Dataset
MSE SNR SSIM
UNet 2.0850e-3 26.8087 0.8930
Conditional DPM 7.1440e-4 30.4605 0.9390
MDA GAN 1.5233e-3 28.1720 0.9063
SeisFusion 4.7471e-4 31.2357 0.9490
V-B Fine-tuning
Although the model trained on the SEG C3 dataset captures some fundamental features of the complex field dataset Mobil Avo Viking Graben Line 12 and exhibits a degree of generalization, the reconstruction performance remains relatively low. Therefore, fine-tuning the model with a smaller amount of data can improve its reconstruction accuracy. To investigate the amount of data required for fine-tuning, we conducted an experiment using 5%, 10%, and 15% of the field data for fine-tuning. For this experiment, 50 traces were continuously missing in Mobil Avo Viking Graben Line 12, with traces set to 0 to represent the missing data. The reconstruction results for models fine-tuned with different data amounts are shown in the Table VII.
TABLE VII: Comparison of Models Fine-tuned With Different Amounts of Data of Four Reconstruction Networks with 50 Consecutive Missing Traces in Field Dataset
MSE SNR SSIM
5% 8.1423e-5 39.8665 0.9832
10% 6.2259e-5 41.8856 0.9942
15% 6.5413e-5 41.6233 0.9948
It can be observed that 5% of the data is sufficient to fine-tune a high-performance model. Increasing the data amount to 10% achieves the model’s optimal state. Therefore, only 10% of the data is needed for fine-tuning to obtain a high-precision reconstruction model.
V-C Interpretability
With the rapid development of deep learning, there is increasing focus on how neural networks operate, and interpretable deep learning models have long been a research goal among scholars. Regarding the working principles of diffusion models, as elaborated in Fig. 1 and the section on Diffusion Model in our article, the neural network in diffusion models is used to estimate the noise scale of data at the current time step $t$. The diffusion model enables theneural network to learn the noise scale corresponding to each time step $t$ through a forward process. During sampling, starting from time step $T$, the neural network estimates the noise scale of $x_{t}$ and samples to obtain $x_{t-1}$, iteratively obtaining $x_{0}$. Regarding how the neural network estimates the noise scale, we extract and plot feature maps of the neural network as well as the noise estimated by the network, as shown in Fig. 12. We selected a case with 50% random missing data reconstruction on the SEG C3 dataset and plotted features extracted by the network at three time steps: $t=5$, $t=50$, and $t=200$. Heatmaps generated from these plots show that when extracting features, the neural network pays more attention to the signal part of the input data. Even at $t=200$ when noise is substantial, signal features can still be extracted. This focused part gradually contributes to denoising to obtain clean data as desired. Based on the extracted feature maps and embedding of the current time step $t$, the decoder of the neural network ultimately outputs an estimate of the noise $\epsilon_{\theta}(x_{t},y,t)$. With $\epsilon_{\theta}(x_{t},y,t)$ obtained, the mean and variance of the entire data distribution can be calculated. Subsequently, data can be sampled based on reparameterization techniques.
V-D Time
In our guided sampling approach, an iterative process is introduced at each sampling step. Although this allows the diffusion model to use the data to guide and constrain the sampling process, it inevitably increases the sampling time, as the entire guided sampling process is longer compared to diffusion models that do not alter the sampling process. To reduce reconstruction time, DIM [41] can be used during reconstruction. The concept of DIM is compatible with our algorithm. When using the DIM algorithm, 50 sampling steps can achieve reconstruction quality equivalent to diffusion model, and fewer sampling steps, such as 20 or even 10 steps, can be used while maintaining acceptable reconstruction quality.
VI Conclusions